# Enabling Regional Explainability by Automatic and Model-agnostic Rule Extraction

### v2

@@ Results
Evaluation criteria
In the rest of this article, we denote the true label of a sample as $y^{*}$, the model’s prediction as $\hat{y}$, the class of interest as the target class $c$, the sample index as $n$, the number of training samples as $N$, and the indicator function as $\mathbb{I}(\cdot)$. Since all the tasks are classification, we also refer to the specified data region of interest as the target region, denoted by $\hat{y}=c$.
Table 1: Rules extracted from different datasets and models by our method AMORE and Decision Tree (DT) classifier. In general, AMORE gives much higher confidence criteria than DT classifier while showing higher fitness even with less support than DT. The results suggest that our method gains a better trade-off between support and confidence for a regional explanation.
@@
Figure 2: Comparing AMORE and DT classifiers while specifying the same minimum support, $s_{min}$, on the $x$ axis and fixing the $l_{max}$ as in Table 1 for each task. We choose the range of $s_{min}$ according to the size of target region for each task. We perform a grid search for other hyperparameters for both methods in all cases and compare the confidence and fitness of their best rule sets. We observe that AMORE generally yields better or equivalent results compared to the DT classifier.
We apply three criteria for evaluating the quality of an extracted rule set $\mathbb{X}$ from a model:
  1. 1.
    $\text{Support}=\sum_{n=1}^{N}\mathbb{I}(x^{(n)}\in\mathbb{X})$
  2. 2.
    Confidence (also called rule accuracy) [11] – the proportion of samples that belong to the target region under the rule set. This metric quantifies the purity of the subspace defined by the rule set. It can be computed as:
  3. 3.
    Fitness – a new metric we propose for regional rule extraction, being the ratio difference between samples that belong to the target region and those that do not, under the rule set, divided by the total number of samples in the target region. The higher the score, the better, and in the best case $\text{Fitness}=1$:
Support and confidence are widely applied in the literature of rule extraction. Typically, support and confidence are negatively correlated because including more samples likely introduces additional noise into the subspace. A better rule set should exhibit reduced correlation by increasing confidence while maintaining a similar support. We define fitness for evaluating regional rule sets, which takes into account the trade-off between support and confidence. This metric measures how well the subspace defined by a rule set fits into the target data region. A rule set with larger support and lower confidence, or one with smaller support and higher confidence, could both result in a lower fitness. Fitness can be negative when the subspace is mostly out of the target region.
Baseline method
@@
In all of our experiments, we chose the Decision Tree (DT) classifier as the baseline method because, similar to our method, it does not require predefined discretisation for numerical features, and we were unable to find another method with this capability. Additionally, it allows us to verify the expected difference between the global and regional rule extraction.
@@
The DT classifiers were trained to classify samples belonging to the target region as positive class and others as negative class. The rules from a DT classifier are formed by the decision path to the node with the highest fitness. The rules from AMORE are the rule set with the highest fitness among all candidate rule sets. All results are provided in Table 1.
@@
Results from AMORE are deterministic when both the model and training set remain unchanged. Similarly, the DT classifier yields consistent results upon varying the random seeds, as all features are involved in split selection in decision trees. Therefore, there is no variance reported in Table 1.
Hyperparameters
@@
In general, there are two hyperparameters required to constrain the extracted rules: i) the maximum number of features included in a rule set is limited by $l_{max}$, constraining the complexity of extracted rules to not be too large; ii) the minimum number of training samples that satisfy the extracted rules is limited by $s_{min}$, constraining the support of the rules to not be too small. We configure the same $l_{max}$ for both methods in each task for consistency. However, the hyperparameter $s_{min}$ is highly correlated to the optimal fitness that a rule set can achieve. Therefore, we conducted a grid search for $s_{min}$, alongside other method-specific hyperparameters, for both methods in each task. We provide the selected hyperparameters for both methods in Table 2 in Supplementary Material. Figure 2 visualises the comparison between the two methods while specifying different values of $s_{min}$.
@@
To avoid selecting rules that have slightly higher fitness but significantly lower confidence, we specify a confidence lower bound $\iota$ for choosing the best rule set across all experiments. This lower bound filters out any rule sets with confidence levels below it when selecting the best rule set based on fitness. However, if no rule set with a confidence above this lower bound is obtained, we simply select the one with the highest fitness. To verify the robustness of AMORE with respect to $\iota$, we conducted a sensitivity analysis by varying $\iota$ from 0.7 to 0.9 (Figure 1 in Supplementary Material). AMORE maintained better or equivalent performance compared to DT classifiers. We chose $\iota=0.8$ as it is considered reasonably good for rule confidence while allowing a higher support.
@@
Table 2: Rules extracted by AMORE for interpreting predictions of individual test samples in the diabetes dataset. We set $s_{min}=1000$, $l_{max}=3$ for all test samples. The prediction threshold $y_{th}$ is 0.457. The rules exhibit lower confidence and fitness when $\hat{y}_{p}$ is lower, indicating that the purity of a local subspace correlates with the model’s uncertainty.
@@
Task 1 – Diabetes prediction
@@
We trained a logistic regression model to classify whether a patient has diabetes or not, and then applied our methods to extract rules that interpret the key knowledge acquired by the model in recognising diabetes patients. We first identify highly influential features using our feature selection method (Algorithm 1) and then extract rules for the target region using those features. The most informative feature identified by both methods is HbA1c level (HbA1c_level). As a measure of a person’s average blood sugar level over the past 2-3 months, it reasonably emerges as a key factor in predicting diabetes, demonstrating the effectiveness of our feature selection method. As only one rule is allowed in this task, AMORE achieved much better performance than the decision tree classifier (Table 1) due to the trade-off issue in global rule extraction is magnified in such scenarios.
Furthermore, we demonstrate obtaining local rules for three test samples by our method, as shown in Table 2: one with a very high predicted probability, another close to the decision threshold, and a third below the threshold. The obtained local rules exhibit lower confidence and fitness when the predicted probability is lower, indicating that the quality of local rules correlates with the model’s uncertainty on the given sample. For samples that are distant from the target region (e.g. $\hat{y}_{p}\approx 0$), our approach may not return valid local rules (rules that give a higher confidence than without them).
Task 2 – Sepsis prediction
@@
The extracted rules can be found in Table 1. The feature Temp indicates temperature (${}^{\circ}\mathrm{C}$) and SPB is Systolic BP (mm Hg). As this task has a high dimensional feature space, our method selected different features with decision trees. We also observe that the top rules of both methods are from features augmented by cumulative time steps. This suggests that the model considers the frequency of taking those vital signs and lab tests as key information for predicting sepsis.
(a) Visualisation of extracted rules from the MNIST dataset. Here, rule 1 is $z_{44}\geq 6.223$ and rule 2 is $0\leq z_{114}\leq 3.105$ as in Table 1.
(b) Visualisation of extracted rules from brain tumour MRI dataset by two samples that satisfy the extracted rules of class “normal”. The above row is a sample truly “normal” and the below row is a sample not labelled as “normal”. The first column highlights pixels that have higher contributions to the latent state ($z_{1200}$) in rule 1 ($z_{1200}\geq 0.653$) and the second row is for the latent state ($z_{1110}$) in rule 2 ($-0.208\leq z_{1110}\leq-0.007$). We visualise the latent state of each rule by colouring pixels according to their impact scores for the latent state, highlighting highly impactful regions of each rule in the figures.
Figure 3: Visualisation of extracted rules for both image datasets.
Task 3 – Molecular toxicity prediction
@@
We trained a graph attention network [31] to predict the categories of compounds using their chemical structures as input. This is a multi-label classification task, and we chose the first two categories for rule extraction as demonstration: the androgen receptor (AR) and aryl hydrocarbon receptor (AhR). We utilise RDKit descriptors [32] to map molecule structures to interpretable features, providing a consistent set of features that describe essential molecule properties. As the results presented in Table 1, AMORE obtained higher fitness than DT classifiers in both cases.
Task 4 –MNIST digits classification
In addition to tabular data, we showcase the capability of our method to interpret models for image data by using MNIST dataset [26]. We trained a neural network with a convolutional layer and two fully connected layers for this task. The target region is set to $\hat{y}=7$, which identifies the model’s key knowledge for classifying the digit 7. We used the model’s latent representations (input to the final linear layer) as features to generate rules, which dimension is 128. The performance of extracted rules can be found in Table 1 as well. For a visual demonstration, we illustrate the extracted rules by samples that satisfy each rule in Figure 3(a). It is clear to see that the first rule emphasises the upper half of digit 7, while the second rule focuses on the lower half.
Task 5 – Brain tumour MRI classification
@@
Like the MNIST task, we also use the latent representations from the model as features to generate rules for a target region in this experiment, which dimension is 1280. We extracted rules for identifying “normal” brain MRI images where the class “normal” is the minority class in this dataset. The sample means of the MRI images are not visually interpretable. Therefore, we showcase two samples that satisfy rules extracted by AMORE in Figure 3(b): one is from class “normal” and the other is malignant. We visualise each rule by colouring pixels according to each pixel’s impact on the latent feature that defines the rule, thereby highlighting the most impactful areas for each rule in an MRI image.
Discussion
@@
In the presented experiments, we illustrate that AMORE can extract representative rules with large support and high confidence for specific target regions, achieving better or equivalent performance compared to DT based explanation. Our proposed approach is also applicable to a broad range of machine learning models without requiring predefined discretisation for numerical features. We compare the best rule sets with only “AND” operations here, however, a more comprehensive view could be constructed by post-processing that merges candidate rule sets using “OR” operations according to customised requirements. We consider this as future work.
@@
While two hyperparameters are configured to limit the minimum support ($s_{min}$) and the maximum length ($l_{max}$) of extracted rules for both methods, we observe that the rules extracted by AMORE often have actual support and length that are closer to these configurations. This aligns with our expectations to have more precise control over regional rule extraction.
@@
We demonstrated how to apply AMORE with data types other than tabular data in several experiments. For instance, in the molecular toxicity prediction task, we map the molecular structures to RDKit descriptors, providing interpretable rules by the descriptors. Likewise, in the MNIST and Brain tumour experiments, we extract rules from the latent space of models and then interpret them through visualisation in the raw pixel space. In general, for data with features that are not directly comprehensible in the form of rules, one can map the raw features to an interpretable feature space so that AMORE can be applicable. For instance, the speech and text data can be interpreted by mapping raw features to semantic concepts, taking into account specific requirements of domain knowledge for an application.

### v3

@@ IV Results
TABLE I: Rules extracted from different datasets and models by our method AMORE and Decision Tree (DT) classifier. In general, AMORE gives much higher confidence criteria than DT classifier while showing higher fitness even with less support than DT. The results suggest that our method gains a better trade-off between support and confidence for a regional explanation.
@@
Figure 5: Comparing AMORE and DT classifiers while specifying the same minimum support, $s_{min}$, on the $x$ axis and fixing the $l_{max}$ as in Table I for each task. We choose the range of $s_{min}$ according to the size of target subgroup for each task. We perform a grid search for other hyperparameters for both methods in all cases and compare the confidence and fitness of their best rule sets. We observe that AMORE generally yields better or equivalent results compared to the DT classifier.
IV-A Baseline method
@@
In all of our experiments, we chose the Decision Tree (DT) classifier as the baseline method because, similar to our method, it is capable of model-agnostic rule extraction and does not require predefined discretization for numerical features. We were unable to find another method with such capabilities. Additionally, it allows us to verify the expected difference between the global and regional rule extraction.
@@
The DT classifiers were trained to classify samples belonging to the target subgroup as positive class and others as negative class. The rules from a DT classifier are formed by the decision path to the node with the highest fitness. The rules from AMORE are the rule set with the highest fitness among all candidate rule sets. All results are provided in Table I.
@@
We set the same random seed for training the classifiers of all tasks. Results from AMORE are deterministic when both the model and training set remain unchanged. Similarly, the DT yields consistent results upon varying the random seeds, as all features are involved in split selection in decision trees. Therefore, there is no variance reported in Table I.
IV-B Hyperparameters
@@
We configure a same $l_{max}$ for both methods in each task. However, the hyperparameter $s_{min}$ is highly correlated to the optimal fitness that a rule set can achieve. Therefore, we conducted a grid search for $s_{min}$, alongside other method-specific hyperparameters, for both methods in each task. We provide the selected hyperparameters for both methods in Table 2 in Supplementary Material. Figure 5 visualizes the comparison between the two methods while varying the value of $s_{min}$.
@@
To avoid selecting rules that have slightly higher fitness but significantly lower confidence, we specify a confidence lower bound $\iota$ for choosing the best rule set during grid search in all experiments. This lower bound filters out any rule sets with confidence levels below it when selecting the best rule set based on fitness. However, if no rule set with a confidence above this lower bound is obtained, we simply select the one with the highest fitness. To verify the robustness of AMORE with respect to $\iota$, we conducted a sensitivity analysis by varying $\iota$ from 0.7 to 0.9 (Figure 1 in Supplementary Material). With different $\iota$, AMORE consistently obtains better or equivalent performance compared to DT classifiers. We chose $\iota=0.8$ as it is considered reasonably good for rule confidence while allowing a higher support.
@@
TABLE II: Rules extracted by AMORE for interpreting predictions of individual test samples with different levels of model uncertainty in the diabetes dataset. We set $s_{min}=1000$, $l_{max}=3$ for all test samples. The prediction threshold $y_{th}$ is 0.457. The rules exhibit lower confidence and fitness when the predicted probability $\hat{y}_{p}$ is lower.
@@
IV-C Task 1 – Diabetes prediction
@@
We trained a logistic regression model to classify whether a patient has diabetes or not, and then extracted rules that interpret the key knowledge acquired by the classifier in recognizing diabetes patients. We first identify highly influential features using our feature selection method (Algorithm 1) and then extract rules using those features. The most informative feature identified by both methods is HbA1c level (HbA1c_level). As a measure of a person’s average blood sugar level over the past 2-3 months, it reasonably emerges as a key factor in predicting diabetes, demonstrating the effectiveness of our feature selection method. As only one rule is allowed in this task, AMORE achieved much better performance than the decision tree classifier (Table I) due to the trade-off issue in global rule extraction is magnified in such scenarios.
Furthermore, we demonstrate obtaining local rules for three test samples by our method, as shown in Table II: one with a very high predicted probability, another close to the decision threshold, and a third below the threshold. The obtained local rules exhibit lower confidence and fitness when the predicted probability is lower, indicating that the quality of local rules correlates with the model’s uncertainty on the given sample. For samples that are distant from the target subgroup (e.g. $\hat{y}_{p}\approx 0$), our approach may not return valid local rules (rules that give a higher confidence than without them).
IV-D Task 2 – Sepsis prediction
@@
As this task has a high dimensional feature space, our method selected different features with decision trees. We also observe that the top rules of both methods are from features augmented by cumulative time steps. This suggests that the model considers the frequency of taking those vital signs and lab tests as key information for predicting sepsis. The extracted rules can be found in Table I. The feature Temp indicates temperature (${}^{\circ}\mathrm{C}$) and SPB is Systolic BP (mm Hg).
Figure 6: Visualization of extracted rules from the MNIST dataset. Here, rule 1 is $z_{44}\geq 6.223$ and rule 2 is $0\leq z_{114}\leq 3.105$ as in Table I.
Figure 7: Visualization of extracted rules from brain tumor MRI dataset by two samples that satisfy the extracted rules of class “normal”. The above row is a sample truly “normal” and the below row is a sample not labeled as “normal”. The first column highlights pixels that have higher contributions to the latent state ($z_{1200}$) in rule 1 ($z_{1200}\geq 0.653$) and the second row is for the latent state ($z_{1110}$) in rule 2 ($-0.208\leq z_{1110}\leq-0.007$). We visualize the latent state of each rule by coloring pixels according to their impact scores for the latent state, highlighting highly impactful regions of each rule in the figures.
IV-E Task 3 – Molecular toxicity prediction
@@
We trained a graph attention network [52] to predict the categories of compounds using their chemical structures as input. This is a multi-label classification task, and we chose the first two categories for rule extraction as demonstration: the androgen receptor (AR) and aryl hydrocarbon receptor (AhR). We utilize RDKit descriptors [53] to map molecule structures to interpretable features, providing a consistent set of features that describe essential molecule properties. As the results presented in Table I, AMORE obtained higher fitness than DT classifiers in both cases.
IV-F Task 4 –MNIST digits classification
In addition to tabular data, we showcase the capability of our method to interpret models for image data by using MNIST dataset [47]. We trained a neural network with a convolutional layer and two fully connected layers for this task. The target class is set to $c=7$, i.e., we identify the model’s key knowledge for classifying the digit 7. We used the model’s latent representations (input to the final linear layer) as features to generate rules, which dimension is 128. The performance of extracted rules can be found in Table I as well. For a visual demonstration, we illustrate the extracted rules by samples that satisfy each rule in Figure 6. It is clear to see that the first rule emphasises the upper half of digit 7, while the second rule focuses on the lower half.
IV-G Task 5 – Brain tumour MRI classification
@@
Like the MNIST task, we also use the latent representations from the model as features to generate rules in this experiment, which dimension is 1280. We extracted rules for identifying “normal” brain MRI images where the class “normal” is the minority class in this dataset. The sample means of the MRI images are not visually interpretable. Therefore, we showcase two samples that satisfy rules extracted by AMORE in Figure 7: one is from class “normal” and the other is malignant. We visualise each rule by colouring pixels according to each pixel’s impact on the latent feature that defines the rule, thereby highlighting the most impactful areas for each rule in an MRI image.
V Discussion
@@
In the presented experiments, we illustrate that AMORE can extract representative rules with large support and high confidence for specific data subgroups, achieving better or equivalent performance compared to DT based explanation. We showcase the best conjunctive rule sets (i.e. with only “AND” operations) here, a more comprehensive view could be constructed by post-processing, for example, using a sequential covering strategy or merging candidate rule sets by “OR” operations.
@@
While two hyperparameters are configured to limit the minimum support ($s_{min}$) and the maximum length ($l_{max}$) of extracted rules for both methods, we observe that the rules extracted by AMORE often have an actual support and length that are closer to these configurations. This aligns with our expectations to have more precise control of extracted rules in regional rule extraction.
@@
We demonstrated how to apply AMORE with data types other than tabular data in several tasks. For instance, in the molecular toxicity prediction task, we map the molecular structures to RDKit descriptors, providing interpretable features by the descriptors. Likewise, in the MNIST and Brain tumor experiments, we extract rules from the latent space of models and then interpret them through visualization in the raw pixel space. In general, for data with features that are not directly comprehensible in the form of logical rules, one can map the raw features to an interpretable feature space so that AMORE can be applicable. For instance, the speech and text data can be transformed by mapping raw features to semantic concepts for interpreting emotion classifiers. However, it is challenging to make such a transformation for time series if we expect interpretable features that can describe complex longitudinal properties. We consider this as future work.